# Model-Based Soft Maximization of Suitable Metrics of Long-Term Human Power

### v1

@@ 1 Introduction
A duty to empower others, especially those with little power, can be defended on consequentialist (Sen 2014), deontological (Hill Jr 2002), and virtue ethical Nussbaum (2019) grounds. At the same time, gradual or sudden disempowerment of humans due to misaligned A(G)I, e.g. due to power-seeking (Turner et al. 2019) or non-corrigible over-optimization of misaligned reward functions (Gao, Schulman, and Hilton 2023), is a key AI safety risk.
Accordingly, some papers explored tasking AI agents explicitly with the empowerment of individual humans, mainly inspired by the information-theoretic channel capacity $\cal E$ between human actions and environmental states (called ‘empowerment’ in (Klyubin, Polani, and Nehaniv 2005)). As that metric is hard to compute, Du et al. (2020) use a proxy based on the variance of terminal states reached under a random policy, while Myers et al. (2024) use contrastive learning of latent state representations to approximate $\cal E$. Salge and Polani (2017) discuss the ‘empowerment’-based approach in more depth for the single-human case.
@@
This paper extends the theoretical foundations of this approach in two ways. We develop an alternative metric of human power that is directly based on the key aspect of power—the ability to attain a wide range of possible goals—and that explicitly and transparently incorporates humans’ knowledge about the state of the world and the actions of the AI agent, their expectations about others’ behavior, e.g. due to social norms, and their bounded rationality. And, using a desiderata-guided approach similar to social choice and welfare theory, we develop an objective function for an AI agent interacting with populations of humans, based on an aggregate of that metric across humans and time that gives the agent desirable incentives such as communicating well (Reddy, Levine, and Dragan 2022), following orders, being corrigible (Potham and Harms 2025), avoiding irreversible changes in the environment, protecting humans and itself from harm and disempowerment, allocating resources fairly and sustainably, and acting “appropriately” by following relevant social norms (Leibo et al. 2024).
@@
Though based on possible human goals, our approach avoids trying to learn individuals’ actual, current goals, because human preferences are changing and non-identifiable (Cao, Cohen, and Szpruch 2021; Banerjee and Duflo 2011) and their prediction unavoidably uncertain (Baker, Saxe, and Tenenbaum 2011). Some critics of a preference-based approach argue for a values-based approach instead (Lowe et al. 2025), which however requires an even more semantic world understanding by the AI agent. A deep semantic understanding is also required in the ‘freedoms’-based conception of AI ethics in (London and Heidari 2024), and compiling its required list of ‘fundamental capabilities’ is difficult (Robeyns 2006). By contrast, like the ‘empowerment’-based approach, our metrics are mostly based on a structural understanding of possibly dynamics, interactions, and transition probabilities and aim to avoid semantic issues by relating human power to the ability to bring about just any possible conditions they might happen to desire.
@@
For theoretical convenience, we work in a model-based setting where the AI agent can plan on the basis of a decent stochastic world model like the “scientist AI” envisioned in Bengio et al. (2025). After developing our human power metric in Section 2, we shortly describe algorithms for softly maximizing it in Section 3, before reporting first insights about the resulting behavior in Section 4.
2 Measuring, Aggregating, and Maximizing Human (Em)Power(ment)
@@
We assume a robot $r$ interacting with several humans $h\in\mathcal{H}$. The robot models the world as a (fully or partially observed) stochastic game with states $s\in\mathcal{S}$ between itself, who gets no extrinsic reward, and the humans who have many possible and potentially changing goals $g_{h}\in\mathcal{G}_{h}$ and get goal-dependent rewards $U_{h}(s^{\prime},g_{h})$ , discounted at factors $\gamma_{h}<1$. Action sets $\mathcal{A}_{r}(s),\mathcal{A}_{h}(s)$ may be state-dependent, action profiles are written $a=(a_{r},a_{\mathcal{H}})=(a_{r},a_{h},a_{-h})$, and the transition kernel is $P(s^{\prime}|s,a)$.
@@
Our task is to design an algorithm for computing a policy $\pi_{r}$ for the robot that implements the aggregate human power maximization objective. For tractability, we formalize this objective as an expected discounted return, $V_{r}=\operatorname*{\mathbb{E}}_{s_{\geq 0}}\sum_{t\geq 0}\gamma_{r}^{t}U_{r}(s_{t})$, based on an intrinsic reward $U_{r}(s)$ representing a suitable assessment of the aggregate human power in $s$. To design $U_{r}(s)$, we first design a power metric $W_{h}(s)$ for individual humans, and then an aggregation function $U_{r}(s)=F_{U}((W_{h}(s))_{h\in\mathcal{H}})$. In designing $W_{h}$ and $F_{U}$, we have several objectives relating to their tractability, interpretability, and the behavioral incentives that they give the robot, in particular relating to AI safety and ethics.
@@
$V_{h}$ has natural interpretation Make $V_{h}$ a goal-reaching probability: $s\in g_{h}$, $s^{\prime}$ reachable from $s\Rightarrow s^{\prime}\notin g_{h}$
@@ 2.1 Individual-Level Metric: ICCEA Power
Our power metric basically measures how many goals a human can achieve. It’s built in three steps: first, defining what it means to achieve a goal; second, adjusting for the human’s own cognitive limits; and third, aggregating this ability across all possible goals into a single number. Rather than trying to capture the full range of subtle aspects of existing notions of human power, we focus on those aspects we believe a robot can robustly infer from the structure of its world model, encoded in state and action sets, transition kernel, and observation functions. Since we want to incentivize the robot to remove constraints and uncertainties, share information, make commitments, and improve human cognition, coordination, and cooperation, our power metric will depend to some degree on $r$’s model of human decision making.
We start with the intuition that humans have ‘goals’ and that ‘power’ is about the means to achieve a wide range of goals. We aim to measure informationally and cognitively constrained effective autonomous (ICCEA) power: essentially how many goals a human can freely choose to reach with more or less certainty, given their information, cognitive capabilities, and others’ behavior. To design the ICCEA power metric, we start with the ability to reach specific goals, then aggregate this over goals, humans, and finally time and uncertainty in ways that have desirable properties.
@@
As an easily interpretable compromise between the reachability (Krakovna et al. 2018) and attainable utility (Turner, Hadfield-Menell, and Tadepalli 2020) approaches, but deviating significantly from the information-theoretic approach (Klyubin, Polani, and Nehaniv 2005) that does not explicitly involve any goals, we assume that each possible temporary goal $g_{h}\in G_{h}$ that $h\in\mathcal{H}$ might have is to reach some set of states, $g_{h}\subseteq\mathcal{S}$, representing a possibly desirable event, and that the corresponding reward function is the indicator function $U_{h}(s,g_{h})=1_{s\in g_{h}}$. In other words, $\mathcal{G}_{h}\subseteq{\cal P}(\mathcal{S})$, the power set of $\mathcal{S}$. We require that the elements of $g_{h}$ are mutually unreachable (e.g., are all terminal states or states for a particular time-point $t$), so that the resulting value $V_{h}^{\pi}(s)$ is simply the probability of the desirable event happening under policy profile $\pi$. This grounds $h$’s goal attainment or ‘attainable utility’ in probability, bounded within [0,1][0,1][ 0 , 1 ], which avoids problems with interpersonal utility comparison such as dominance by “utility monsters”. To avoid problems around probability zero, we also require that $G_{h}$ is wide enough so that each possible state trajectory makes some of $h$’s possible goals fulfilled. [^1] Otherwise $X_{h}^{-\xi}$ in eqn. (10) could get infinite. We believe that restricting the model to this very basic type of goal will simplify the derivation of $G_{h}$ from learned latent representations of generic world states and generic human goals as encoded in language or foundation models, will obviate the need for individual-level data about a particular human’s possible goals, and will thus mitigate risks arising from misaligned models of possible human goals. [^2] An even more restrictive choice would be to identify goals with single states, but this seem implausibly specific in complex, partially observed, multi-agent environments, and it would lead to $G_{h}$ not behaving well under world model refinements.
@@
[…]
Then $r$ assumes $h$ uses a mixture (governed by parameter $\nu_{h}$) between habitual, ‘system-1’-type behavior encoded in some default policy $\pi^{0}_{h}(s,g_{h})$, again reflecting social norms, and boundedly rational, ‘system-2’-type behavior represented by a (Boltzmann) softmax policy with rationality parameter $\beta_{h}$ (which $r$ might estimate from observations, Safari et al. (2024)), see eq. (4). As in Ghosal et al. (2023), $\beta_{h}$ may be state-dependent, which will give $r$ incentives to choose states with larger $\beta_{h}$, see below. Since they reflect social norms, $\nu_{h}$, $\mu_{-h}$, and $\pi^{0}_{h}$ are the only significantly “semantically loaded” elements of the world model.
Regarding $r$’s own actions, $r$ models $s$ as containing information about what actions $r$ has committed publicly to choose from, so that $A_{r}(s)$ only contains those actions, and then $r$ models $h$ as being maximally cautious regarding $r$’s remaining actions, hence using the $\min_{a_{r}\in\mathcal{A}_{r}(s)}$ operator in (3). The reason for this is not realism but that it incentivizes $r$ to make goal-independent binding commitments about its interaction behavior, e.g. by properly labelling its buttons or promising to react to certain verbal commands in certain ways. The incentive for $r$ to choose such a “commitment action” in some state $s$ arises since that will make $\mathcal{A}_{r}(s^{\prime})$ of later states $s^{\prime}$ smaller and will thus weakly increase $Q^{m}_{h}(s^{\prime},g_{h},a_{h})$, without $r$ ever having to know $g_{h}$!
These assumptions also allow $r$ to first calculate $\pi_{h}$ for each $h$ independently and before deciding on $\pi_{r}$, thus avoiding potential issues around non-uniqueness of strategic equilibria or non-stationarity in learning, see ‘Phase 1’ below.
@@
While $r$ assumes $h$’s behavior $\pi_{h}$ is based on $h$’s own value assessment $V^{m}_{h}$, $r$’s assessment of $h$’s effective goal-reaching ability, $V^{e}_{h}$, that $r$ will use in computing $h$’s effective power, will generally differ from $V^{m}_{h}$. This is because $r$’s beliefs about other humans’ policies, $\pi_{-h}$, may differ from $h$’s beliefs $\mu_{-h}$, and $r$’s policy $\pi_{r}$ will generally differ from $h$’s assumption $\min_{a_{r}}$. This is reflected in eq. (8), which calculates the probability of $g_{h}$ being fulfilled if $r$ and $h$ use $\pi_{r}$ and $\pi_{h}$, averaged over others’ potential goals $g_{-h}$ and corresponding behaviors $\pi_{-h}$.
@@
How should $r$ aggregate $h$’s effective goal attainment abilities $V^{e}_{h}(s,g_{h})$ across all possible goals $g_{h}\in\mathcal{G}_{h}$ to get an assessment of $h$’s ICCEA power in $s$, $W_{h}(s)$? We require the aggregation to be independent of goal labels, continuous, strictly increasing in each $V^{e}_{h}(s,g_{h})$, and to fulfil another natural axiom (“independence of unaffected goals”). Similar to Fleming (1952), this implies that the aggregation must be “separable”, $W_{h}(s)=F^{G}(X_{h}(s))$ with $X_{h}(s)=\sum_{g_{h}}f^{G}(V^{e}_{h}(s,g_{h}))$, for some continuous and strictly increasing functions $F^{G},f^{G}$. Nicely, since $X_{h}(s)=\operatorname*{\mathbb{E}}_{g_{h}}|G_{h}|f^{G}(V^{e}_{h}(s,g_{h}))$, we can then hope to use stochastic approximation to learn it (see below).
We choose the inner transformation $f^{G}$ to be of the form $f^{G}(v)=v^{\zeta}$ for some $\zeta>0$ for reasons that will become clear when discussing $r$’s reward function below. This is also analogous to certain “non-expected” utility theories, in particular rank-dependent utility theory with a probability-weighting function $w(p)=p^{\zeta}$ (Quiggin 1982). By choosing $\zeta>1$, we incentivize $r$ to reduce uncertainty, because it would then consider $h$ to be more powerful when $h$ can choose between two deterministic outcomes (so that $X_{h}(s)=2\times 1^{\zeta}=2$) than when $h$ can choose between two coin tosses (so that $X_{h}(s)=4\times(1/2)^{\zeta}<2$). This can be interpreted as risk aversion or a preference for reliability.
Because the subsequent aggregation across humans will involve an additional transformation anyway, our choice of the outer transformation $F^{G}$ is somewhat arbitrary. We choose $F^{G}=\log_{2}$ so that power is measured in bits and, at least in the limit of full rationality, it behaves in an additive way if the game is decomposable into several independent simultaneous games. This choice also leads to a convenient relationship between our final ICCEA power metric,
@@
[…]
@@
Klyubin, Polani, and Nehaniv (2005) define ‘empowerment’ as the channel capacity between actions and states. In a single-player multi-armed bandit environment with possible outcomes $s^{\prime}$, this is the maximal mutual information $E_{h}=\textstyle\max_{\pi_{h}}\mathbb{I}_{\pi_{h}}(a_{h};s^{\prime})$. In the Supplement, we show that $E_{h}\leq W_{h}$ if we put $G_{h}=\mathcal{S}$, $\zeta=1$, and assume full rationality ($\nu_{h}=0$, $\beta_{h}=\infty$), similar to what Myers et al. (2024) have shown. Similarly, for $\zeta>1$, our $W_{h}$ is an upper bound of an entropy-regularized version of ‘empowerment’,
@@
sharing the same value range and coinciding in edge cases. Note that while the policy $\pi_{h}$ that $r$ estimates in our approach is a function of state and goal $g_{h}$, has typically low entropy as it aims to reach $g_{h}$, and can be found by standard dynamic programming or RL approaches, the maximizing “policy” $\pi_{h}$ in (2) has no natural interpretation, has typically high entropy to maximize $\mathbb{I}(a_{h};s^{\prime})$, and is harder to find since the optimization problem is non-convex.

### v2

@@ 1 Introduction
A duty to empower others, especially those with little power, can be defended on consequentialist (Sen 2014), deontological (Hill Jr 2002), and virtue ethical (Nussbaum 2019) grounds. At the same time, gradual or sudden disempowerment of humans due to misaligned A(G)I, e.g. due to power-seeking (Turner et al. 2019) or non-corrigible over-optimization of misaligned reward functions (Gao, Schulman, and Hilton 2023), is a key AI safety risk.
[…]
@@
This paper extends the theoretical foundations of this approach in two ways. First, we develop an alternative metric of human power, ‘ICCEA power’, that is directly based on the key aspect of power: the ability to attain a wide range of possible goals. Our metric explicitly and transparently incorporates humans’ knowledge about the actions of the AI agent, their expectations about others’ behavior, e.g. due to social norms, and their bounded rationality. Second, using an approach guided by desiderata similar to the axioms of social choice and welfare theory, we develop an objective function for an AI agent interacting with populations of humans, based on an aggregate of ICCEA power across humans and time that gives the agent desirable incentives such as communicating well (Reddy, Levine, and Dragan 2022), following orders, being corrigible (Potham and Harms 2025), avoiding irreversible changes in the environment, protecting humans and itself from harm and disempowerment, allocating resources fairly and sustainably, and acting “appropriately” by following relevant social norms (Leibo et al. 2024).
@@
Though based on possible human goals, our approach avoids trying to learn individuals’ actual, current goals, because human preferences are changing and non-identifiable (Cao, Cohen, and Szpruch 2021; Banerjee and Duflo 2011) and their prediction is unavoidably uncertain (Baker, Saxe, and Tenenbaum 2011). Some critics of a preference-based approach argue for a values-based approach instead (Lowe et al. 2025), which however requires an even more semantic world understanding by the AI agent. A deep semantic understanding is also required in the ‘freedoms’-based conception of AI ethics in (London and Heidari 2024), and compiling its required list of ‘fundamental capabilities’ is difficult (Robeyns 2006). By contrast, like the ‘empowerment’-based approach, our metrics are mostly based on a structural understanding of possibly dynamics, interactions, and transition probabilities and aim to avoid semantic issues by relating human power to the ability to bring about just any possible conditions a human might happen to desire.
@@
For theoretical convenience, we work in a model-based setting where the AI agent can plan on the basis of a decent stochastic world model like the “scientist AI” envisioned in Bengio et al. (2025). After developing our human power metric in Section 2, we shortly describe algorithms for softly maximizing it in Section 3, before reporting insights about the resulting behavior in Section 4. Section 5 concludes.
2 Measuring, Aggregating, and Softly Maximizing Human (Em)Power(ment)
@@
We assume a robot $r$ interacting with several humans $h\in\mathcal{H}$. The robot models the world as a (fully or partially observed) stochastic game with states $s\in\mathcal{S}$. It gets no extrinsic reward itself. Humans have many possible and potentially changing goals $g_{h}\in\mathcal{G}_{h}$ and get goal-dependent rewards $U_{h}(s^{\prime},g_{h})$, discounted at factors $\gamma_{h}<1$. Action sets $\mathcal{A}_{r}(s),\mathcal{A}_{h}(s)$ may be state-dependent. Action and policy profiles are written $a=(a_{r},a_{\mathcal{H}})=(a_{r},a_{h},a_{-h})$, $\pi=(\pi_{r},\pi_{\mathcal{H}})$. The transition kernel is $P(s^{\prime}|s,a)$.
@@
Task
We want to design an algorithm for computing a policy $\pi_{r}$ for the robot that implements the aggregate human power maximization objective. For tractability, we formalize this objective as an expected discounted return, $V_{r}=\operatorname*{\mathbb{E}}_{s_{\geq 0}}\sum_{t\geq 0}\gamma_{r}^{t}U_{r}(s_{t})$, based on an intrinsic reward $U_{r}(s)$ which represents a suitable assessment of the aggregate human power in $s$. To design $U_{r}(s)$, we first design a power metric $W_{h}(s)$ for individual humans, and then an aggregation function $U_{r}(s)=F_{U}((W_{h}(s))_{h\in\mathcal{H}})$.
In designing $W_{h}$ and $F_{U}$, we have several objectives relating to their tractability, interpretability, and the behavioral incentives that they give the robot.
$~~{}\times\text{$\beta_{h}(s,g_{h})$-softmax for~}Q^{m}_{h}(s,g_{h},\cdot),$ (2)
$\pi_{r}(s)(a)$ $\propto(-Q_{r}(s,a_{r}))^{-\beta_{r}},$ (5)
Table 1: Computation of ICCEA power $W_{h}=\log_{2}X_{h}$ and intrinsic robot reward $U_{r}$ as derived in the text.
@@
$V_{h}$ has natural interpretation Make $V_{h}$ a goal-reaching probability: $\in g_{h}, ss^{\prime}$ reachable from $s\Rightarrow s^{\prime}\notin g_{h}$
@@
Our final equations for the fully observed case are collected in Table 1 (see the Supplement for the partially observed case). We will motivate them now in detail.
@@ 2.1 Individual-Level Metric: ICCEA Power
Our power metric basically counts the effective number of goals a human can achieve. It is built in three steps: defining what it means to achieve a goal, adjusting for human bounded rationality, and aggregating the achievement ability across all possible goals into one number. Rather than trying to capture the full range of subtle aspects of existing notions of human power, we focus on those aspects we believe a robot can robustly infer from the structure of its world model, encoded in state and action sets, transition kernel, and observation functions. Since we want to incentivize the robot to remove constraints and uncertainties, share information, make commitments, and improve human cognition, coordination, and cooperation, our power metric will also depend on $r$’s model of human decision making.
We start with the intuition that humans have ‘goals’ and that ‘power’ is about the means to achieve a wide range of goals. We aim to measure informationally and cognitively constrained effective autonomous (ICCEA) power: essentially how many goals a human can freely choose to reach with more or less certainty, given their information, cognitive capabilities, and others’ behavior.
@@
We use an easily interpretable compromise between the reachability (Krakovna et al. 2018) and attainable utility (Turner, Hadfield-Menell, and Tadepalli 2020) approaches. We deviate significantly from the information-theoretic approach (Klyubin, Polani, and Nehaniv 2005) which does not explicitly involve any goals. We assume that each possible temporary goal that $h\in\mathcal{H}$ might have is to reach some set of states $g_{h}\subseteq\mathcal{S}$ representing a possibly desirable event. The corresponding reward function is the indicator function $U_{h}(s,g_{h})=1_{s\in g_{h}}$. So the set of goals $\mathcal{G}_{h}$ is simply a set of subsets of $\mathcal{S}$. We require that the desirable states in a goal $g_{h}$ are mutually unreachable (e.g., are all terminal states or states for a particular time-point $t$). Then the resulting value $V_{h}^{\pi}(s)$ is simply the probability of the desirable event under policy profile $\pi$. This grounds $h$’s goal attainment or ‘attainable utility’ in probability and increases interpretability. As it is then bounded within $[0,1]$, this also avoids aggregation problems such as dominance by “utility monsters”. To avoid issues with probability zero, $\mathcal{G}_{h}$ must be wide enough so that each possible state trajectory fulfills at least one possible goal. [^1] Otherwise the quantity $X_{h}^{-\xi}$ in eqn. (8) could get infinite. E.g., $\mathcal{G}_{h}$ could be a partition of the terminal states.
We believe that restricting the model to this very basic type of goal will simplify the derivation of $\mathcal{G}_{h}$ from learned latent representations of generic world states and generic human goals as encoded in language or foundation models, and will obviate the need for individual-level data about a particular human’s possible goals. This should mitigate risks arising from misaligned models of human goals. [^2] An even more restrictive choice would be to identify goals with single (terminal) states, but this seems too specific in complex, partially observed, multi-agent environments, and the resulting $\mathcal{G}_{h}$ would not behave well under world model refinements.
@@
We equip $r$ with a simple model of $h$’s decision making that focuses on giving $r$ the right incentives. It assumes $h$ cannot realize the maximal goal attainment probability due to a variety of reasons relating to exploration, imperfect action implementation, information constraints, others’ behavior, and potentially state-dependent cognitive limitations. In particular, $r$ does not assume $h$ to have correct beliefs about others’ behavior that would lead to an equilibrium (Nash, quantal response, etc.). Instead, $r$ models $h$ as having fixed beliefs $\mu_{-h}$ about other humans’ behavior (where ‘h-h’ is short for $\mathcal{H}\setminus\{h\}$). These would also reflect social norms (which LLM-based systems already understand, Smith et al. (2024)). Hence the state-goal-action values $Q^{m}_{h}(s,g_{h},a_{h})$ that $r$ assumes guide $h$’s behavior are based on $\mu_{-h}$. This is reflected in eq. (1) in Table 1.
Regarding human actions, $r$ assumes that $h$ uses a mixture (governed by a probability $\nu_{h}$) between (i) habitual, ‘system-1’ behavior encoded in some default policy $\pi^{0}_{h}(s,g_{h})$, again reflecting social norms, and (ii) boundedly rational, ‘system-2’ behavior represented by a Boltzmann policy with rationality parameter $\beta_{h}$, see eq. (2). As in Ghosal et al. (2023), $\beta_{h}$ may be state-dependent, which will give $r$ incentives to choose states with larger $\beta_{h}$, and it might be estimated from observations (Safari et al. 2024). Reflecting social norms, $\nu_{h}$, $\mu_{-h}$, and $\pi^{0}_{h}$ are the only significantly “semantically loaded” elements of the world model.
To model human beliefs about the robot’s actions, the world model contains information about what actions $r$ has previously committed to choose from in a state $s$: the action set $\mathcal{A}_{r}(s)$ only contains those actions, and different commitment histories are considered different states. Then $r$ models $h$ as being cautious regarding $r$’s commitment-compliant actions. It thus uses the $\min_{a_{r}\in\mathcal{A}_{r}(s)}$ operator to compute $Q^{m}_{h}(s,g_{h},a_{h})$ in eq. (1). This is not for realism but because it incentivizes $r$ to make goal-independent commitments about its interaction behavior, e.g. by properly labelling its buttons or promising to react to certain verbal commands in certain ways. The incentive for $r$ to choose such a “commitment action” in some state $s$ arises since that will make $\mathcal{A}_{r}(s^{\prime})$ of later states $s^{\prime}$ smaller and will thus weakly increase $Q^{m}_{h}(s^{\prime},g_{h},a_{h})$, without $r$ ever having to guess $g_{h}$.
These assumptions also allow $r$ to first calculate a behavior prior $\pi_{h}$ for each $h$ independently, before deciding its own policy $\pi_{r}$. This avoids issues around non-uniqueness of strategic equilibria and non-stationarity in learning.
@@
While $r$ assumes $h$’s behavior $\pi_{h}$ is based on $h$’s cautious value assessment $V^{m}_{h}$, eq. (3), its assessment of $h$’s effective goal-reaching ability $V^{e}_{h}$, used in computing $h$’s power, will generally differ from $V^{m}_{h}$. This is because $r$’s beliefs about other humans’ policies, $\pi_{-h}$, may differ from $h$’s beliefs, $\mu_{-h}$. And $r$’s policy $\pi_{r}$ will generally differ from the worst case given by $\min_{a_{r}}$. Hence eq. (6) calculates $V^{e}_{h}$ as the probability of $g_{h}$ being fulfilled under the actual policy $\pi_{r}$ and the derived prior $\pi_{\mathcal{H}}$, averaged over others’ potential goals $g_{-h}$ and corresponding behaviors $\pi_{-h}$. This is then the basis of the power metric.
@@
How should $r$ aggregate $h$’s effective goal attainment abilities $V^{e}_{h}(s,g_{h})$ across all possible goals $g_{h}\in\mathcal{G}_{h}$ to get an assessment of $h$’s ICCEA power in $s$, $W_{h}(s)$? Similar to Fleming (1952), we want the aggregation to be independent of goal labels, continuous, strictly increasing in each $V^{e}_{h}(s,g_{h})$, and to fulfil “independence of unaffected goals”. This implies that the aggregation must be “separable”, $W_{h}(s)=F^{G}(X_{h}(s))$ with $X_{h}(s)=\sum_{g_{h}}f^{G}(V^{e}_{h}(s,g_{h}))$ for some continuous and strictly increasing transformations $F^{G},f^{G}$. Since $X_{h}(s)=\operatorname*{\mathbb{E}}_{g_{h}}|G_{h}|f^{G}(V^{e}_{h}(s,g_{h}))$, we can then also hope to learn it via stochastic approximation.
For reasons that will become clear when discussing $r$’s reward function $U_{r}$ below, we choose the inner transformation $f^{G}$ to be of the form $f^{G}(v)=v^{\zeta}$ for some $\zeta>0$. This is analogous to certain “non-expected” utility theories, in particular to rank-dependent utility theory with the probability-weighting function $w(p)=p^{\zeta}$ (Quiggin 1982). We choose $\zeta>1$ to incentivize $r$ to reduce uncertainty. It would then consider $h$ to be more powerful when $h$ can choose between two deterministic outcomes (so that $X_{h}(s)=2\times 1^{\zeta}=2$) than when $h$ can choose between two coin tosses (so that $X_{h}(s)=4\times(1/2)^{\zeta}<2$). This can be interpreted as risk aversion or a preference for reliability.
Our choice of the outer transformation $F^{G}$ is somewhat arbitrary. This is because the subsequent aggregation across humans will involve an additional transformation anyway. We choose $F^{G}=\log_{2}$ so that power is measured in bits. In the limit of full rationality, it then also behaves in an additive way if the game is decomposable into several independent simultaneous games. This choice leads to a convenient relationship between our final ICCEA power metric,
@@
and the information-theoretic notion of ‘empowerment’, as described below. In the case where $h$ can choose between fulfilling $k$ different goals for sure, we simply have $W_{h}(s)=\log_{2}k$. Because each trajectory fulfils at least one goal $g_{h}$, we always have $X_{h}(s)>0$ and thus $W_{h}(s)>$-$\infty$. But $W_{h}(s)$ may be negative in situations with very little control.333 If $\mathcal{G}_{h}$ is a partition of $\mathcal{S}$ into $k$ blocks, the world is deterministic, and $h$ cannot influence it, then $W_{h}(s)=0$. If instead each $g_{h}$ is reached with probability $1/k$ regardless of what $h$ does, and if $\zeta=2$, then $W_{h}(s)$ attains its minimum of $\log_{2}k$. The resulting symmetric value range $W_{h}(s)\in[-\log_{2}k,\log_{2}k]$ suggests making $\mathcal{G}_{h}$ a partition and putting $\zeta=2$ might be a natural choice.
@@
Klyubin, Polani, and Nehaniv (2005) define ‘empowerment’ as the channel capacity between actions and states. In a single-player multi-armed bandit environment with possible outcomes $s^{\prime}$, this equals the maximal mutual information $E_{h}=\textstyle\max_{\pi_{h}}\mathbb{I}_{\pi_{h}}(a_{h};s^{\prime})$. In the Supplement, we show that $E_{h}\leq W_{h}$ if we put $\mathcal{G}_{h}=\mathcal{S}$, $\zeta=1$, and assume full rationality ($\nu_{h}=0$, $\beta_{h}=\infty$), similar to what Myers et al. (2024) have shown. Similarly, for $\zeta>1$, our metric $W_{h}$ is an upper bound of an entropy-regularized version of ‘empowerment’,
@@
sharing the same value range and coinciding in edge cases.
But while the policy $\pi_{h}$ that $r$ estimates in our approach is a function of state and goal $g_{h}$, has typically low entropy as it aims to reach $g_{h}$, and can be found by standard dynamic programming or RL approaches, the maximizing “policy” $\pi_{h}$ in eq. (11) has no real use, has typically high entropy in order to maximize $\mathbb{I}(a_{h};s^{\prime})$, and is harder to find since the optimization problem is non-convex.